# Zeyu Zhu*, Kevin Qinghong Lin*, Mike Zheng Shou🖂

### v1

@@ Abstract
Academic presentation videos have become an essential medium for research communication, yet producing them remains highly labor-intensive, often requiring hours of slide design, recording, and editing for a short 2 to 10 minutes video. Unlike natural video, presentation video generation involves distinctive challenges: long-context inputs from research papers, dense multi-modal information (text, figures, tables), and the need to coordinate multiple aligned channels such as slides, subtitles, speech, and human talker. To address these challenges, we introduce Paper2Video, the first benchmark of 101 research papers paired with author-created presentation videos, slides, and speaker metadata. We further design four tailored evaluation metrics—Meta Similarity, PresentArena, PresentQuiz, and IP Memory—to measure how videos convey the paper’s information to the audience. Building on this foundation, we propose PaperTalker, the first multi-agent framework for academic presentation video generation. It integrates slide generation with effective layout refinement by a novel effective Tree Search Visual Choice, cursor grounding, subtitling, speech synthesis, and talking-head rendering, while parallelizing slide-wise generation for efficiency. Experiments on Paper2Video demonstrate that the presentation videos produced by our approach are more faithful and informative than existing baselines, establishing a practical step toward automated and ready-to-use academic video generation. Our dataset, agent, and code are available at https://github.com/showlab/Paper2Video
@@ 3.3 Evaluation Metrics
IP Memory – How videos affect the author’s visibility and work impact? Another key purpose of academic presentation videos is to enhance the visibility and impact of the author’s work. Yet, this metric is unclear and difficult to simulate and thus remains an open problem. In real-conference settings, audiences who recall a scholar after attending their presentation are more inclined to pose relevant questions in later interactions. Motivated by this phenomenon, we propose a metric to assess how effectively a presentation video enables the audience to recall the work. Additional implementation details are provided in Appendix B.1.
@@ 5.2 Main Results
Table 2: Detailed evaluation result of Paper2Video across three baselines. $\text{PaperTalk}^{*}$ represents a simple version without presenter and cursor. Bold and Underline indicates the best and the second.
@@ 6 Conclusions
Appendix A Checklist
A.1  The Use of Large Language Models
In our work, LLMs are used for following aspects:
  • Using an LLM to help with paper writing. We use GPT5 to help optimize language, correct grammar and write table code.
  • Using an LLM as a research assistant. We use GPT5 to help search related works.
  • Using an LLM in our methods and experiment. This is described in the paper.
A.2  Ethics Statement
We confirm that our study did not use any sensitive data where all data are public available. We have conducted this research and reported our findings responsibly. All results are presented transparently, including both performance gains and any observed limitations. We have diligently cited all relevant prior work and data sources to give proper credit and context. By following best practices in documentation and research integrity, we aim to contribute positively to the scientific community while upholding the highest ethical standards.
A.3  Reproducibility statement
We are committed to ensuring the reproducibility of our results. All code and data needed to reproduce the experiments will be made publicly available. We will release this repository openly with an appropriate open-source license upon publication. The datasets used in our experiments are standard public benchmarks for language modeling and understanding (e.g., widely-used corpora and evaluation sets). These resources are readily accessible to other researchers.
Appendix B Evaluation Metrics
B.1  IP Memory
@@
Appendix C Experiment
C.1  Video Results
@@
C.2  Results of Tree Search Visual Choice

### v2

# Zeyu Zhu* Kevin Qinghong Lin* Mike Zheng Shou🖂

@@ Abstract
Academic presentation videos have become an essential medium for research communication, yet producing them remains highly labor-intensive, often requiring hours of slide design, recording, and editing for a short 2 to 10 minutes video. Unlike natural video, presentation video generation involves distinctive challenges: long-context inputs from research papers, dense multi-modal information (text, figures, tables), and the need to coordinate multiple aligned channels such as slides, subtitles, speech, and human talker. To address these challenges, we introduce Paper2Video, the first benchmark of 101 research papers paired with author-created presentation videos, slides, and speaker metadata. We further design four tailored evaluation metrics—Meta Similarity, PresentArena, PresentQuiz, and IP Memory—to measure how videos convey the paper’s information to the audience and affect the work impact. Building on this foundation, we propose PaperTalker, the first multi-agent framework for academic presentation video generation. It integrates slide generation with effective layout refinement by a novel effective Tree Search Visual Choice, cursor grounding, subtitling, speech synthesis, and talking-head rendering, while parallelizing slide-wise generation for efficiency. Experiments on Paper2Video demonstrate that the presentation videos produced by our approach are more faithful and informative than existing baselines, establishing a practical step toward automated and ready-to-use academic video generation. Our dataset, agent, and code are available at https://github.com/showlab/Paper2Video.
@@ 3.3 Evaluation Metrics
IP Memory – How videos affect the author’s visibility and work impact? Another key purpose of academic presentation videos is to enhance the visibility and impact of the author’s work. Yet, this metric is unclear and difficult to simulate and thus remains an open problem. In real-conference settings, audiences who recall a scholar after attending their presentation are more inclined to pose relevant questions in later interactions. Motivated by this phenomenon, we propose a metric to assess how effectively a presentation video enables the audience to recall the work. Additional implementation details are provided in Appendix A.1.
@@ 5.2 Main Results
Efficiency Analysis. As shown in Table 3, PaperTalker achieves the lowest cost. This efficiency stems from our slide-generation design: adopting Beamer reduces token usage for slide creation, and our tree search visual choice module which is a lightweight post-processing step. Runtime is further reduced by our parallel talking-head generation mechanism. By contrast, PresentAgent Shi et al. [2025] incurs higher token costs due to frequent refinement queries during slide editing.
Table 2: Detailed evaluation result of Paper2Video across three baselines. $\text{PaperTalk}^{*}$ represents a simple version without presenter and cursor. Bold and Underline indicates the best and the second. NA means the results of method are not applicable to the metric.
@@ 6 Conclusions
Appendix A Evaluation Metrics
A.1  IP Memory
@@
Appendix B Experiment
B.1  Video Results
@@
B.2  Results of Tree Search Visual Choice